OpenStudy (anonymous):

Which of the following functions is a solution to the differential equation y" + 2y' + y = 0 possible answers: y = et y = -e t y = e -t y = e -2t i tryed solving it put i only get : y(x) = c1e^(-x)+c2e^(-x) x

OpenStudy (anonymous):

\[y''+2y'+y=0\] gives the characteristic equation \[r^2+2r+1=(r+1)^2=0\] which has root \(r=-1\) of multiplicity 2, which means the solution would be of the form \[y=C_1e^{-x}+C_2xe^{-x}\] By the principle of superposition, either one of these terms is considered a solution, so \(y=e^{-x}\) is a solution if \(C=1\).